Calculus II, Spring 2009 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 11

Assigned May 07, 2009.

1. Section 14.6: Problems 13, 15, 17, 19, 25, 33, 41, 43.

2. Chap 14: 28, 37, 39, 41 (need not evaluate).